Calculate the elastic potential energy when a 1000N/m spring is stretched by a 20N force
1
Expert's answer
2019-11-01T11:24:43-0400

The Hooke's law says


"F=kx"

So, the deformation of a spring


"x=\frac{F}{k}=\frac{20\:\rm N}{1000\:\rm N/m}=0.02\:\rm m"

The elastic potential energy


"E_p=\frac{kx^2}{2}=\frac{1000\:\rm N/m\times (0.02\:\rm m)^2}{2}=0.2\:\rm J"